# Retrieval-Guided Compositional Image Generation for Cross-Domain Few-Shot Object Detection

### v1

@@ Abstract
Cross-Domain Few-Shot Object Detection (CD-FSOD) aims to detect novel objects with only a handful of labeled samples from previously unseen domains. While data augmentation and generative methods have shown promise in few-shot learning, their effectiveness for CD-FSOD remains unclear due to the need for both visual realism and domain alignment. Existing strategies, such as copy-paste augmentation and text-to-image generation, often fail to preserve the correct object category or produce backgrounds coherent with the target domain, making them non-trivial to apply directly to CD-FSOD. To address these challenges, we propose Domain-RAG, a training-free, retrieval-guided compositional image generation framework tailored for CD-FSOD. Domain-RAG consists of three stages: domain-aware background retrieval, domain-guided background generation, and foreground-background composition. Specifically, the input image is first decomposed into foreground and background regions. We then retrieve semantically and stylistically similar images to guide a generative model in synthesizing a new background, conditioned on both the original and retrieved contexts. Finally, the preserved foreground is composed with the newly generated domain-aligned background to form the generated image. Without requiring any additional supervision or training, Domain-RAG produces high-quality, domain-consistent samples across diverse tasks, including CD-FSOD, remote sensing FSOD, and camouflaged FSOD. Extensive experiments show consistent improvements over strong baselines and establish new state-of-the-art results. Codes will be released upon acceptance.
@@ 1 Introduction
Cross-Domain Few-Shot Object Detection (CD-FSOD) [1], an emerging task derived from cross-domain few-shot learning (CD-FSL) [2], aims to tackle few-shot object detection (FSOD) across different domains. Unlike conventional FSOD [3], which assumes source and target data share similar distributions, CD-FSOD considers more realistic scenarios with significant domain shifts, for example, transferring from natural images to industrial anomaly images, remote sensing imagery, or underwater environments. By simultaneously involving the challenges of few-shot learning and domain shift, CD-FSOD poses significant challenges for existing detectors.
Due to the extreme scarcity of labeled data, e.g., as few as 1 or 5 annotated samples per category, a natural and intuitive solution is to leverage data augmentation to alleviate the data bottleneck. Although image augmentation and generation techniques have been extensively studied and shown effective in other few-shot learning tasks [4, 5, 6, 7], it remains unclear whether they can produce high-quality training samples for CD-FSOD. Different from the few-shot classification or in-domain FSOD, this setting requires not only accurate object annotation but also strong domain consistency, which has not been tackled in prior data augmentation-based few-shot learning methods.
@@
To generate training images for CD-FSOD, the most straightforward approach is copy-paste (Fig. 1(a)). While easy to implement, such images often lack realism and domain coherence. A more advanced strategy is to build on recent generative models, particularly the trending text-to-image generation, such as SDXL [8], FLUX [9]. Most existing methods in this area focus on synthesizing foreground objects via text prompts, such as "a photo of category" (Fig. 1(b)). However, they might fail to preserve the object semantics when applied to novel categories and domains. Such a category shift is problematic for CD-FSOD, which has to tackle fine-grained objects and the domain gap. Other approaches generate diverse backgrounds (Fig. 1(c)), guided by text descriptions of the image. While this better preserves the foreground, purely textual descriptions often fall short of capturing precise domain characteristics and struggle to ensure semantic and visual consistency between foreground and background. These limitations motivate us to develop a new image generation framework capable of synthesizing visually coherent, domain-aligned training samples for CD-FSOD. Specifically, we aim to: ① preserve the original foreground object, ② generate diverse backgrounds that are both semantically and stylistically aligned with the query image and its domain, and ③ produce visually realistic images with valid annotations suitable for downstream detection training.
@@
To that end, we propose Domain-RAG, a retrieval-guided compositional image generation framework built upon the principle of fix the foreground, adapt the background. Leveraging the nature of object detection, Domain-RAG begins by decomposing the target image into its foreground object and background, where the background is recovered by applying an inpainting model [10] to the object-masked region. Although simple in principle, this step is critical for preserving the original object and its annotations, laying the foundation for controllable compositional generation. The core challenge then lies in generating a new background that is semantically and stylistically compatible with the foreground. Inspired by the paradigm of retrieval-augmented generation (RAG) [11], we inject structured visual priors into the generative process to guide background synthesis. As illustrated in Fig. 1(e), Domain-RAG consists of the following three stages: 1) Domain-Aware Background Retrieval. We introduce an image database (e.g., COCO [12]) containing diverse natural scenes, from which we retrieve candidate backgrounds that are semantically and stylistically similar to the inpainted background of the target image. Semantic similarity is computed using high-level visual features, while style similarity is measured via style-based descriptors [13]. 2) Domain-Guided Background Generation. Rather than using the retrieved backgrounds directly, we feed them along with the target’s inpainted background into a generative model to synthesize a new background that better reflects the visual characteristics of the target domain. To ensure compatibility with modern diffusion models, Redux [14] is applied to convert visual image cues into descriptive text prompts, enabling direct use of text-to-image generation models. 3) Foreground-Background Composition. Finally, the preserved foreground is seamlessly composed onto the synthesized, domain-aligned background using a mask-guided generative model. The resulting image maintains the original object while embedding it in a realistic, domain-consistent context (Fig. 1(d)). The entire Domain-RAG pipeline is training-free and can be directly integrated with existing detectors without any additional supervision or retraining, making it particularly suitable for low-shot scenarios such as 1-shot CD-FSOD.
@@ 2 Related Works
Cross-Domain Few-Shot Tasks. Few-shot learning across domains has been widely studied [2, 15, 16, 6, 17, 18, 19, 20, 21, 22, 23, 24, 25], but most works focus only on classification. The more realistic task of cross-domain few-shot object detection (CD-FSOD) [1, 26], which involves both recognizing and localizing objects, remains underexplored. Recent methods like CD-ViTO [1] and ETS [27] address CD-FSOD. CD-ViTO introduces the task with a closed-source setting (COCO as the only source), while ETS uses a more practical open-source setting [26] and leverages data augmentation via pretrained GroundingDINO [28]. In this paper, we adopt the open-source setting and further improve augmentation using retrieval-guided compositional generation.
@@
FSOD Beyond Domains. Beyond classic CD-FSOD tasks, many FSOD or detection problems also involve domain shifts [29, 30, 31, 32], even if not explicitly labeled as cross-domain. Two notable examples are Remote Sensing FSOD (RS-FSOD) [33] and Camouflaged FSOD [34]. RS-FSOD uses remote sensing images, which differ from natural scenes in color, perspective, and resolution, creating a clear domain gap. Camouflaged FSOD involves detecting objects that blend into their backgrounds—like fish underwater or animals in the wild—posing challenges for generalization. We include both tasks to assess our method under diverse and difficult cross-domain scenarios.
@@
Data Augmentation. It is a key technique for improving object detection. Traditional methods like copy-paste [35], cropping, and color jittering [36] are simple but offer limited semantic variety. Recently, generative models—especially text-to-image models like ControlNet [37], SDXL [8], FLUX [9], and FLUX-Fill [38] have enabled more advanced augmentations. Methods such as X-Paste [39], Lin et al. [7], and Zhang et al. [40] generate new foregrounds to paste on diverse backgrounds, while others [41, 42, 37, 43, 44] use text prompts to jointly create foregrounds and backgrounds. However, these methods typically rely on large amounts of in-domain data for training, which limits their adaptability to novel categories or unseen domains. In contrast, our Domain-RAG is training-free and leverages retrieved real-world images as visual priors to generate domain-consistent samples, making it well-suited for CD-FSOD.
@@
Retrieval-Augmented Generation in Vision. First introduced in NLP [11], retrieval-augmented generation (RAG) enhances outputs by incorporating relevant retrieved content as external knowledge. Its strong performance has led to applications in vision tasks such as image captioning [45, 46], visual question answering [47, 48], and image generation [49, 50, 51]. However, current RAG-based image generation methods are aimed at open-ended synthesis and are not suited for object detection, particularly in cross-domain few-shot settings, where both domain alignment and object fidelity are crucial. To the best of our knowledge, we are the first to introduce a RAG-inspired, training-free image generation framework specifically designed for CD-FSOD.
@@ 3 Proposed Method
Problem Setup. The CD-FSOD task aims to adapt an object detector from a source domain $\mathcal{D}_{S}$ to a target domain $\mathcal{D}_{T}$, where the data distributions $\mathcal{P}_{S}$ and $\mathcal{P}_{T}$ differ. We use the few-shot setting, i.e., $N$-way $K$-shot protocol to evaluate detection results in $\mathcal{D}_{T}$. Specifically, a support set $\mathcal{S^{\textit{N}\times\textit{K}}}\subset\mathcal{D}_{T}$ provides $K$ labeled examples per novel class, and a query set $\mathcal{Q}$ is used for evaluation. We use the open-source setting introduced in the 1st CD-FSOD Challenge [26], which allows foundation models pretrained on large-scale data, to explore the potential of foundation models in CD-FSOD. Particularly, instead of pretraining on $\mathcal{D}_{S}$, we directly finetune a pretrained detector (e.g., GroundingDINO [28]) on the support set $\mathcal{S}$ and evaluate the results on the query set $\mathcal{Q}$. To mitigate the limited size of $\mathcal{S}$, we augment each support instance with $n$ synthetic images, effectively expanding each class from $K$ to $K\times(n+1)$ examples.
Overview. We propose Domain-RAG—a novel, training-free, retrieval-guided compositional generation framework that enhances support diversity by generating domain-aligned samples. To enable retrieval, we use COCO [12] as the database $\mathcal{D}_{base}$, serving as a gallery of candidate backgrounds. Following the core principle of "fix the foreground, adapt the background", Domain-RAG processes each support image $x\in\mathcal{S}$ by first decomposing it into foreground object(s) and background. As shown in Fig. 2, the framework then proceeds through three key stages: 1) domain-aware background retrieval first obatins the inpainted background $b_{init}$ from $x$ and then retrieves $G$ candidate backgrounds $b_{re}$ from $\mathcal{D}_{base}$ that are semantically and stylistically similar. 2) domain-guided background generation feeds each $\{b_{init},b_{re}\}$ pair into a generative model to synthesize a new domain-aligned background $b_{dom}$. 3) foreground-background composition finally produces $n$ new images $x^{+}$ by compositing the preserved foreground onto each $b_{dom}$ using a mask-guided generative model.
Figure 2: Illustration of our Domain-RAG. Built on our principle of "fix foreground, adapt background", we first decompose image and process it with three key modules: domain-aware background retrieval, domain-guided background generation, and foreground-background composition.
@@ 3.1 Domain-Aware Background Retrieval
In practice, given a support image $x$, we remove the ground-truth bounding box with LaMa inpainting [10] to obtain a background without foreground $b_{init}$. We use the CLIP encoder to extract embeddings from the initial background $b_{init}$ and the database $\mathcal{D}_{base}$, which we refer to as $F_{bg}$ and $F_{base}$, respectively. We compute cosine similarity between the visual feature of the current background query $F_{bg}$ and the CLIP embedding of each sample in the database. The top $m$ most similar images are selected based on this similarity ranking, forming the candidate set $B_{<clip,m>}$, which contains $m$ images of the form $b_{clip}$. The subscript notation indicates that the set is constructed using CLIP vision envoder and contains $m$ elements.
@@ 3.2 Domain-Guided Background Generation
To fully leverage the retrieved images while keeping the generation process training-free, we adopt the Flux-Redux model[14] to encode each image into a prompt embedding. Given our domain-aware retrieval results $B_{<re,\ n>}$, let $\text{redux}(\cdot)$ denote FLUX-Redux encoder and $\text{FLUX}(\cdot)$ denote the FLUX generator. For each support image, we extract its clean background embedding $F_{bg}=\text{redux}(b_{init})$ and the embedding of the top retrieved image $b_{re}\in B_{<re,\ n>}$ as $F_{re}=\text{redux}(b_{re})$. We then fuse them as $F_{dom}=\lambda_{1}F_{bg}+\lambda_{2}F_{re}$, where $\lambda_{1}$ and $\lambda_{2}$ are hyper parameters.
@@ 4 Experiments
Setups. We conduct experiments on three FSOD tasks with domain shifts: 1) CD-FSOD: Following the CD-ViTO benchmark [1], we evaluate on six diverse target domains: ArTaxOr [52] (photorealistic), Clipart1k [53] (cartoon), DIOR [54] (aerial), DeepFish [55] (underwater), NEU-DET [56] (industrial), and UODD [57] (underwater). 2) Remote Sensing FSOD (RS-FSOD): In addition to DIOR, we include NWPU VHR-10 [58], a popular remote sensing dataset for FSOD. 3) Camouflaged FSOD: We also test on CAMOS-FS [34], a recent dataset with 47 categories where objects are deliberately camouflaged into the background. For each task, we follow the standard dataset splits and evaluation protocols: 1/5/10-shot for CD-FSOD, 3/5/10/20-shot for RS-FSOD, and 1/2/3/5-shot for Camouflaged FSOD. Results are reported using mean Average Precision (mAP).
Implementation Details. We use pretrained GroundingDINO [28] with Swin-Transformer [59] Base (Swin-B) as backbone as our baseline. For the retrieval stage, the hyper parameters are set to $m=100$, $n=5$, throughout all experiments. For the background generation stage, fusion hyper parameters $\lambda_{1}$ and $\lambda_{2}$ are set to 1.0 and 0.8 respectively. We fine-tune the model for $30$ epochs by default, but reduce to $5$ for faster-converging datasets like Clipart1k and DeepFish. We use AdamW [60] with learning rate and weight decay set to $1\times 10^{-4}$, and we scale the backbone’s learning rate by $0.1$. All experiments are run on four Tesla V100 GPUs or eight 5880 Ada GPUs, or a single A800 GPU. Further details are in the Appendix.
@@
1-shot Meta-RCNN  [61] ResNet50 2.8 - 7.8 - - 3.6 /
TFA w/cos  [62] ResNet50 3.1 - 8.0 - - 4.4 /
FSCE  [63] ResNet50 3.7 - 8.6 - - 3.9 /
DeFRCN  [64] ResNet50 3.6 - 9.3 - - 4.5 /
Distill-cdfsod  [65] ResNet50 5.1 7.6 10.5 nan nan 5.9 /
ViTDeT-FT  [66] ViT-B/14 5.9 6.1 12.9 0.9 2.4 4.0 5.4
Detic [67] ViT-L/14 0.6 11.4 0.1 0.9 0.0 0.0 2.2
Detic-FT  [67] ViT-L/14 3.2 15.1 4.1 9.0 3.8 4.2 6.6
DE-ViT [68] ViT-L/14 0.4 0.5 2.7 0.4 0.4 1.5 1.0
CD-ViTO [1] ViT-L/14 21.0 17.7 17.8 20.3 3.6 3.1 13.9
@@
5-shot Meta-RCNN  [61] ResNet50 8.5 - 17.7 - - 8.8 /
TFA w/cos  [62] ResNet50 8.8 - 18.1 - - 8.7 /
FSCE  [63] ResNet50 10.2 - 18.7 - - 9.6 /
DeFRCN  [64] ResNet50 9.9 - 18.9 - - 9.9 /
Distill-cdfsod  [65] ResNet50 12.5 23.3 19.1 15.5 16.0 12.2 16.4
ViTDeT-FT  [66] ViT-B/14 20.9 23.3 23.3 9.0 13.5 11.1 16.9
Detic [67] ViT-L/14 0.6 11.4 0.1 0.9 0.0 0.0 2.2
Detic-FT  [67] ViT-L/14 8.7 20.2 12.1 14.3 14.1 10.4 13.3
DE-ViT [68] ViT-L/14 10.1 5.5 7.8 2.5 1.5 3.1 5.1
CD-ViTO [1] ViT-L/14 47.9 41.1 26.9 22.3 11.4 6.8 26.1
@@
10-shot Meta-RCNN  [61] ResNet50 14.0 - 20.6 - - 11.2 /
TFA w/cos  [62] ResNet50 14.8 - 20.5 - - 11.8 /
FSCE  [63] ResNet50 15.9 - 21.9 - - 12.0 /
DeFRCN  [64] ResNet50 15.5 - 22.9 - - 12.1 /
Distill-cdfsod  [65] ResNet50 18.1 27.3 26.5 15.5 21.1 14.5 20.5
ViTDeT-FT  [66] ViT-B/14 23.4 25.6 29.4 6.5 15.8 15.6 19.4
Detic [67] ViT-L/14 0.6 11.4 0.1 0.9 0.0 0.0 2.2
Detic-FT  [67] ViT-L/14 12.0 22.3 15.4 17.9 16.8 14.4 16.5
DE-ViT [68] ViT-L/14 9.2 11.0 8.4 2.1 1.8 3.1 5.9
CD-ViTO [1] ViT-L/14 60.5 44.3 30.8 22.3 12.8 7.0 29.6

### v2

@@ Abstract
Cross-Domain Few-Shot Object Detection (CD-FSOD) aims to detect novel objects with only a handful of labeled samples from previously unseen domains. While data augmentation and generative methods have shown promise in few-shot learning, their effectiveness for CD-FSOD remains unclear due to the need for both visual realism and domain alignment. Existing strategies, such as copy-paste augmentation and text-to-image generation, often fail to preserve the correct object category or produce backgrounds coherent with the target domain, making them non-trivial to apply directly to CD-FSOD. To address these challenges, we propose Domain-RAG, a training-free, retrieval-guided compositional image generation framework tailored for CD-FSOD. Domain-RAG consists of three stages: domain-aware background retrieval, domain-guided background generation, and foreground-background composition. Specifically, the input image is first decomposed into foreground and background regions. We then retrieve semantically and stylistically similar images to guide a generative model in synthesizing a new background, conditioned on both the original and retrieved contexts. Finally, the preserved foreground is composed with the newly generated domain-aligned background to form the generated image. Without requiring any additional supervision or training, Domain-RAG produces high-quality, domain-consistent samples across diverse tasks, including CD-FSOD, remote sensing FSOD, and camouflaged FSOD. Extensive experiments show consistent improvements over strong baselines and establish new state-of-the-art results. The source code and instructions are available at https://github.com/LiYu0524/Domain-RAG.
@@ 1 Introduction
Cross-Domain Few-Shot Object Detection (CD-FSOD) [fu2024cross], an emerging task derived from cross-domain few-shot learning (CD-FSL) [guo2020broader], aims to tackle few-shot object detection (FSOD) across different domains. Unlike conventional FSOD [kohler2023few], which assumes source and target data share similar distributions, CD-FSOD considers more realistic scenarios with significant domain shifts, for example, transferring from natural images to industrial anomaly images, remote sensing imagery, or underwater environments. By simultaneously involving the challenges of few-shot learning and domain shift, CD-FSOD poses significant challenges for existing detectors.
Due to the extreme scarcity of labeled data, e.g., as few as 1 or 5 annotated samples per category, a natural and intuitive solution is to leverage data augmentation to alleviate the data bottleneck. Although image augmentation and generation techniques have been extensively studied and shown effective in other few-shot learning tasks [zhou2023revisiting, li2023knowledge, liang2021boosting, lin2023explore], it remains unclear whether they can produce high-quality training samples for CD-FSOD. Different from the few-shot classification or in-domain FSOD, this setting requires not only accurate object annotation but also strong domain consistency, which has not been tackled in prior data augmentation-based few-shot learning methods.
@@
To generate training images for CD-FSOD, the most straightforward approach is copy-paste (Fig. 1(a)). While easy to implement, such images often lack realism and domain coherence. A more advanced strategy is to build on recent generative models, particularly the trending text-to-image generation, such as SDXL [podell2023sdxl], FLUX [FLUX]. Most existing methods in this area focus on synthesizing foreground objects via text prompts, such as "a photo of category" (Fig. 1(b)). However, they might fail to preserve the object semantics when applied to novel categories and domains. Such a category shift is problematic for CD-FSOD, which has to tackle fine-grained objects and the domain gap. Other approaches generate diverse backgrounds (Fig. 1(c)), guided by text descriptions of the image. While this better preserves the foreground, purely textual descriptions often fall short of capturing precise domain characteristics and struggle to ensure semantic and visual consistency between foreground and background. These limitations motivate us to develop a new image generation framework capable of synthesizing visually coherent, domain-aligned training samples for CD-FSOD. Specifically, we aim to: ① preserve the original foreground object, ② generate diverse backgrounds that are both semantically and stylistically aligned with the query image and its domain, and ③ produce visually realistic images with valid annotations suitable for downstream detection training.
@@
To that end, we propose Domain-RAG, a retrieval-guided compositional image generation framework built upon the principle of fix the foreground, adapt the background. Leveraging the nature of object detection, Domain-RAG begins by decomposing the target image into its foreground object and background, where the background is recovered by applying an inpainting model [suvorov2022resolution] to the object-masked region. Although simple in principle, this step is critical for preserving the original object and its annotations, laying the foundation for controllable compositional generation. The core challenge then lies in generating a new background that is semantically and stylistically compatible with the foreground. Inspired by the paradigm of retrieval-augmented generation (RAG) [lewis2020retrieval], we inject structured visual priors into the generative process to guide background synthesis. As illustrated in Fig. 1(e), Domain-RAG consists of the following three stages: 1) Domain-Aware Background Retrieval. We introduce an image database (e.g., COCO [lin2014microsoft]) containing diverse natural scenes, from which we retrieve candidate backgrounds that are semantically and stylistically similar to the inpainted background of the target image. Semantic similarity is computed using high-level visual features, while style similarity is measured via style-based descriptors [huang2017arbitrary]. 2) Domain-Guided Background Generation. Rather than using the retrieved backgrounds directly, we feed them along with the target’s inpainted background into a generative model to synthesize a new background that better reflects the visual characteristics of the target domain. To ensure compatibility with modern diffusion models, Redux [fluxredux2024] is applied to convert visual image cues into descriptive text prompts, enabling direct use of text-to-image generation models. 3) Foreground-Background Composition. Finally, the preserved foreground is seamlessly composed onto the synthesized, domain-aligned background using a mask-guided generative model. The resulting image maintains the original object while embedding it in a realistic, domain-consistent context (Fig. 1(d)). The entire Domain-RAG pipeline is training-free and can be directly integrated with existing detectors without any additional supervision or retraining, making it particularly suitable for low-shot scenarios such as 1-shot CD-FSOD.
@@ 2 Related Works
Cross-Domain Few-Shot Tasks. Few-shot learning across domains has been widely studied [guo2020broader, tseng2020cross, fu2021meta, liang2021boosting, fu2022wave, zhuo2022tgdm, fu2022me, zhang2022free, fu2023styleadv, luo2023closer, zhuo2024prompt, zou2024closer, peng2024advancing, zhuo2024unified], but most works focus only on classification. The more realistic task of cross-domain few-shot object detection (CD-FSOD) [fu2024cross, fu2025ntire], which involves both recognizing and localizing objects, remains underexplored. Recent methods like CD-ViTO [fu2024cross] and ETS [pan2025enhance] address CD-FSOD. CD-ViTO introduces the task with a closed-source setting (COCO as the only source), while ETS uses a more practical open-source setting [fu2025ntire] and leverages data augmentation via pretrained GroundingDINO [liu2024grounding]. In this paper, we adopt the open-source setting and further improve augmentation using retrieval-guided compositional generation.
@@
FSOD Beyond Domains. Beyond classic CD-FSOD tasks, many FSOD or detection problems also involve domain shifts [cao2024chasing, lu2025musia, lu2025enhancing, liu2024multi, an2025generalized, liu2025ot, anmultimodality, pan2025locate], even if not explicitly labeled as cross-domain. Two notable examples are Remote Sensing FSOD (RS-FSOD) [liu2024few] and Camouflaged FSOD [nguyen2024art]. RS-FSOD uses remote sensing images, which differ from natural scenes in color, perspective, and resolution, creating a clear domain gap. Camouflaged FSOD involves detecting objects that blend into their backgrounds—like fish underwater or animals in the wild—posing challenges for generalization. We include both tasks to assess our method under diverse and difficult cross-domain scenarios.
@@
Data Augmentation. Data augmentation [wen2025rohoi, peng2024mitigating] is a key technique for the vision community. Traditional methods for object detection, like copy-paste [ghiasi2021simple], cropping, and color jittering [bochkovskiy2020yolov4], are simple but offer limited semantic variety. Recently, generative models—especially text-to-image models like ControlNet [zhang2023adding], SDXL [podell2023sdxl], FLUX [FLUX], and FLUX-Fill [fluxfill2024] have enabled more advanced augmentations. Methods such as X-Paste [zhao2023x], Lin et al. [lin2023explore], and Zhang et al. [zhang2024advancing] generate new foregrounds to paste on diverse backgrounds, while others [ni2022imaginarynet, zhang2023diffusionengine, zhang2023adding, liu2023spatio, chen2023geodiffusion, zhou2023seeds, pan2025earthsynth, liu2025control] use text prompts to jointly create foregrounds and backgrounds. However, these methods typically rely on large amounts of in-domain data for training, which limits their adaptability to novel categories or unseen domains. In contrast, our Domain-RAG is training-free and leverages retrieved real-world images as visual priors to generate domain-consistent samples, making it well-suited for CD-FSOD.
@@
Retrieval-Augmented Generation in Vision. First introduced in NLP [lewis2020retrieval], retrieval-augmented generation (RAG) enhances outputs by incorporating relevant retrieved content as external knowledge. Its strong performance has led to applications in vision tasks such as image captioning [ramos2023smallcap, li2024evcap], visual question answering [lin2023fine, he2024g], and image generation [blattmann2022retrieval, lyu2025realrag, zheng2025retrieval], and pose estimation [wang2025rag]. However, current RAG-based image generation methods are aimed at open-ended synthesis and are not suited for object detection, particularly in cross-domain few-shot settings, where both domain alignment and object fidelity are crucial. To the best of our knowledge, we are the first to introduce a RAG-inspired, training-free image generation framework specifically designed for CD-FSOD.
@@ 3 Proposed Method
Problem Setup. The CD-FSOD task aims to adapt an object detector from a source domain $\mathcal{D}_{S}$ to a target domain $\mathcal{D}_{T}$, where the data distributions $\mathcal{P}_{S}$ and $\mathcal{P}_{T}$ differ. We use the few-shot setting, i.e., $N$-way $K$-shot protocol to evaluate detection results in $\mathcal{D}_{T}$. Specifically, a support set $\mathcal{S^{\textit{N}\times\textit{K}}}\subset\mathcal{D}_{T}$ provides $K$ labeled examples per novel class, and a query set $\mathcal{Q}$ is used for evaluation. We use the open-source setting introduced in the 1st CD-FSOD Challenge [fu2025ntire], which allows foundation models pretrained on large-scale data, to explore the potential of foundation models in CD-FSOD. Particularly, instead of pretraining on $\mathcal{D}_{S}$, we directly finetune a pretrained detector (e.g., GroundingDINO [liu2024grounding]) on the support set $\mathcal{S}$ and evaluate the results on the query set $\mathcal{Q}$. To mitigate the limited size of $\mathcal{S}$, we augment each support instance with $n$ synthetic images, effectively expanding each class from $K$ to $K\times(n+1)$ examples.
Overview. We propose Domain-RAG—a novel, training-free, retrieval-guided compositional generation framework that enhances support diversity by generating domain-aligned samples. To enable retrieval, we use COCO [lin2014microsoft] as the database $\mathcal{D}_{base}$, serving as a gallery of candidate backgrounds. Following the core principle of "fix the foreground, adapt the background", Domain-RAG processes each support image $x\in\mathcal{S}$ by first decomposing it into foreground object(s) and background. As shown in Fig. 2, the framework then proceeds through three key stages: 1) domain-aware background retrieval first obtains the inpainted background $b_{init}$ from $x$ and then retrieves $G$ candidate backgrounds $b_{re}$ from $\mathcal{D}_{base}$ that are semantically and stylistically similar. 2) domain-guided background generation feeds each $\{b_{init},b_{re}\}$ pair into a generative model to synthesize a new domain-aligned background $b_{dom}$. 3) foreground-background composition finally produces $n$ new images $x^{+}$ by compositing the preserved foreground onto each $b_{dom}$ using a mask-guided generative model.
Figure 2: Illustration of our Domain-RAG. Built on our principle of "fix the foreground, adapt the background", we first decompose image and process it with three key modules: domain-aware background retrieval, domain-guided background generation, and foreground-background composition.
@@ 3.1 Domain-Aware Background Retrieval
In practice, given a support image $x$, we remove the ground-truth bounding box with LaMa inpainting [suvorov2022resolution] to obtain a background without foreground $b_{init}$. We use the CLIP encoder to extract embeddings from the initial background $b_{init}$ and the database $\mathcal{D}_{base}$, which we refer to as $F_{bg}$ and $F_{base}$, respectively. We compute cosine similarity between the visual feature of the current background query $F_{bg}$ and the CLIP embedding of each sample in the database. The top $m$ most similar images are selected based on this similarity ranking, forming the candidate set $B_{<clip,m>}$, which contains $m$ images of the form $b_{clip}$. The subscript notation indicates that the set is constructed using CLIP vision encoder and contains $m$ elements.
@@ 3.2 Domain-Guided Background Generation
To fully leverage the retrieved images while keeping the generation process training-free, we adopt the Flux-Redux model[fluxredux2024] to encode each image into a prompt embedding. Given our domain-aware retrieval results $B_{<re,\ n>}$, let $\text{redux}(\cdot)$ denote FLUX-Redux encoder and $\text{FLUX}(\cdot)$ denote the FLUX generator. For each support image, we extract its clean background embedding $F_{bg}=\text{redux}(b_{init})$ and the embedding of the top retrieved image $b_{re}\in B_{<re,\ n>}$ as $F_{re}=\text{redux}(b_{re})$. We then fuse them as $F_{dom}=\lambda_{1}F_{bg}+\lambda_{2}F_{re}$, where $\lambda_{1}$ and $\lambda_{2}$ are hyper parameters.
@@ 4 Experiments
Setups. We conduct experiments on three FSOD tasks with domain shifts: 1) CD-FSOD: Following the CD-ViTO benchmark [fu2024cross], we evaluate on six diverse target domains: ArTaxOr [GeirArTaxOr] (photorealistic), Clipart1k [inoue2018cross] (cartoon), DIOR [li2020object] (aerial), DeepFish [saleh2020realistic] (underwater), NEU-DET [song2013noise] (industrial), and UODD [jiang2021underwater] (underwater). 2) Remote Sensing FSOD (RS-FSOD): In addition to DIOR, we include NWPU VHR-10 [niemeyer2014contextual], a popular remote sensing dataset for FSOD. 3) Camouflaged FSOD: We also test on CAMO-FS [nguyen2024art], a recent dataset with 47 categories where objects are deliberately camouflaged into the background. For each task, we follow the standard dataset splits and evaluation protocols: 1/5/10-shot for CD-FSOD, 3/5/10/20-shot for RS-FSOD, and 1/2/3/5-shot for Camouflaged FSOD. Results are reported using mean Average Precision (mAP).
@@
1-shot
Meta-RCNN  [yan2019meta] ResNet50 2.8 - 7.8 - - 3.6 /
TFA w/cos  [wang2020frustratingly] ResNet50 3.1 - 8.0 - - 4.4 /
FSCE  [sun2021fsce] ResNet50 3.7 - 8.6 - - 3.9 /
DeFRCN  [qiao2021defrcn] ResNet50 3.6 - 9.3 - - 4.5 /
Distill-cdfsod  [xiong2023cd] ResNet50 5.1 7.6 10.5 nan nan 5.9 /
ViTDeT-FT  [li2022exploring] ViT-B/14 5.9 6.1 12.9 0.9 2.4 4.0 5.4
Detic [zhou2022detecting] ViT-L/14 0.6 11.4 0.1 0.9 0.0 0.0 2.2
Detic-FT  [zhou2022detecting] ViT-L/14 3.2 15.1 4.1 9.0 3.8 4.2 6.6
DE-ViT [zhang2023detect] ViT-L/14 0.4 0.5 2.7 0.4 0.4 1.5 1.0
CD-ViTO [fu2024cross] ViT-L/14 21.0 17.7 17.8 20.3 3.6 3.1 13.9
@@
5-shot
Meta-RCNN  [yan2019meta] ResNet50 8.5 - 17.7 - - 8.8 /
TFA w/cos  [wang2020frustratingly] ResNet50 8.8 - 18.1 - - 8.7 /
FSCE  [sun2021fsce] ResNet50 10.2 - 18.7 - - 9.6 /
DeFRCN  [qiao2021defrcn] ResNet50 9.9 - 18.9 - - 9.9 /
Distill-cdfsod  [xiong2023cd] ResNet50 12.5 23.3 19.1 15.5 16.0 12.2 16.4
ViTDeT-FT  [li2022exploring] ViT-B/14 20.9 23.3 23.3 9.0 13.5 11.1 16.9
Detic [zhou2022detecting] ViT-L/14 0.6 11.4 0.1 0.9 0.0 0.0 2.2
Detic-FT  [zhou2022detecting] ViT-L/14 8.7 20.2 12.1 14.3 14.1 10.4 13.3
DE-ViT [zhang2023detect] ViT-L/14 10.1 5.5 7.8 2.5 1.5 3.1 5.1
CD-ViTO [fu2024cross] ViT-L/14 47.9 41.1 26.9 22.3 11.4 6.8 26.1
@@
10-shot
Meta-RCNN  [yan2019meta] ResNet50 14.0 - 20.6 - - 11.2 /
TFA w/cos  [wang2020frustratingly] ResNet50 14.8 - 20.5 - - 11.8 /
FSCE  [sun2021fsce] ResNet50 15.9 - 21.9 - - 12.0 /
DeFRCN  [qiao2021defrcn] ResNet50 15.5 - 22.9 - - 12.1 /
Distill-cdfsod  [xiong2023cd] ResNet50 18.1 27.3 26.5 15.5 21.1 14.5 20.5
ViTDeT-FT  [li2022exploring] ViT-B/14 23.4 25.6 29.4 6.5 15.8 15.6 19.4
Detic [zhou2022detecting] ViT-L/14 0.6 11.4 0.1 0.9 0.0 0.0 2.2
Detic-FT  [zhou2022detecting] ViT-L/14 12.0 22.3 15.4 17.9 16.8 14.4 16.5
DE-ViT [zhang2023detect] ViT-L/14 9.2 11.0 8.4 2.1 1.8 3.1 5.9
CD-ViTO [fu2024cross] ViT-L/14 60.5 44.3 30.8 22.3 12.8 7.0 29.6
@@
132
Implementation Details. We use pretrained GroundingDINO [liu2024grounding] with Swin-Transformer [liu2021swin] Base (Swin-B) as backbone as our baseline. For the retrieval stage, the hyper parameters are set to $m=100$, $n=5$, throughout all experiments. For the background generation stage, fusion hyper parameters $\lambda_{1}$ and $\lambda_{2}$ are set to 1.0 and 0.8 respectively. We fine-tune the model for $30$ epochs by default, but reduce to $5$ for faster-converging datasets like Clipart1k and DeepFish. We use AdamW [loshchilov2017decoupled] with learning rate and weight decay set to $1\times 10^{-4}$, and we scale the backbone’s learning rate by $0.1$. All experiments are run on four Tesla V100 GPUs or eight 5880 Ada GPUs, or a single A800 GPU. Further details are in the Appendix.